The numbers 3, 4, and 5 are called Pythagorean triples since 3 to the second power + 4 to second power = 5 to the second power. The numbers 5, 12, and 13 are also Pythagorean triples since 5 to the second power + 12 to the second power = 13 to the second power. Can you find at least 5 more Pythagorean triples? Actually, there is a set of formulas that will generate an infinite number of Pythagorean triples. What are they? I need to write a brief report on the subject.
Answer by mananth(16946)

Pythagorean triples are very common in algebra and geometry.
8,6,10
5,12, 13
15, 8 , 17
12, 16 , 20
7 24,25
16,30 34
Let m & n be two positive integers n> m.
Then n^2-m^2 , 2mn , n^2 + m^2 is a pythagorean triple
You can assign valus for n & m and check it up.
suppose you take 7 &6
n^2-m^2= 7^-6^= 13
2mn = 84
n^2+m^2= 85
13^2 + 84^2 = 85^2
cheers
Ananth